Elimination method
6x-y=27
7y=6x-81

Found 2 solutions by stanbon, jim_thompson5910:
Answer by stanbon(75887)

Elimination method
6x-y=27
7y=6x-81
-----
Rearrange::
6x - y = 27
6x - 7y = 81
--------
Subtract and solve for "y"::
6y = -54
y = -9
------------
Solve for "x" using 6x-y = 27
6x + 9 = 27
6x = 18
x = 3
------------
Cheers,
Stan H.
------------

Answer by jim_thompson5910(35256)

I'll get you started


6x-y=27
7y=6x-81


6x-y=27
-6x+7y=-81


Now add the equations


6x-y=27
-6x+7y=-81
-----------
0x+6y=-54

6y = -54 leads to y = -9

Use this to find x.